1. find w if 67.1=29.7-0.2w.
2. Betty is thinking of 2 consecutive integers whose sum is 41. Let x represent the smaller unknown integer.
a. How could you represent the larger unknown integer in terms of x?
b. Write an equation showing that the sum of the two unknown integers is 41.
c. Solve your equation. What integers is Betty thinking of?

Answer by AdolphousC(70)

1.
67.1 = 29.7 - 0.2W Subtract 29.7 from both sides
37.4 = -0.2W divide both sides by -0.2
-187 = W


a. Consecutive integers are integers that are one after the other, so if X is the smaller of two integers, then the next integer is simply X + 1. ( if you have 5 how do you get to the next larger integer by arithmatic? Add 1 )


b. our two integers are X and X + 1, add them together and set equal to 41
X + ( X + 1 ) = 41


c. Solve by first combining like terms
2X + 1 = 41 Subtract 1 on both sides
2X = 40 Divide by 2 on both sides
X = 20
If X = 20, then 20 is your smaller integer, so the larger integer X + 1 is 20 + 1 or 21.
Betty's two numbers are 20 and 21